10 ml of N/2 KOH, 30 ml of N/10 NaOH, 75 ml of N/5 Mg(OH)2 are mixed together. The normality of resulting solution is?
A) .02N
B) .2N
C) 0.5N
D) 0.05 N

Grade:11

2 Answers

Shashank Shekhar
13 years ago

Total Milliequivalents=(10*0.5)+(30*0.1)+(75*0.2)=23

Total Volume=10+30+75=115 ml

Normality=23/115=0.2N

Note---If Normality of Mg(OH)2 is N then normality of Hydroxide ion is also N and not 2N.

           But if molarity of Mg(OH)2 is M then molarity of hydroxide ion is 2M.

DHANARAJ
7 years ago
HEI THERETHE NORMALITY OF FINAL SOLUTION IS DETERMINED BY THE CONCENTRATION OF OH- INFACT ,SO CALCULATING THAT IT COMES OUT TO BE 23 mmol =23m